Ticket Number: HEA-NOC/20050426-5 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20050426 15:59 UTC+1 Problem Start: 20050427 00:00 UTC+1
Ticket Closed: 20050427 08:53 UTC+1 Problem End: 20050427 00:20 UTC+1
Due to a hardware issue on the INEX router the router will be rebooted at
midnight. Traffic will reroute through other peering arrangements.
Cable & Wireless, Colt Internet,NTL,INEX Route Collector BT Ireland (ipv4
only),Eircom Net - HEAnet (will switch to backup connection on sw03)
20050426 15:46 SY There will be a slight preformance hit as traffic
20050427 08:51 SY INEX peering restored to normal after 5 minutes, work
completed successfully, traffic restored to normal.
This ticket can be monitored at http://www.hea.net/tickets/20050426-5